Multiple Choice
Suppose P(A) = 0.60,P(B) = 0.85,and A and B are independent.The probability of the complement of the event (A and B) is:
A) .4 * .15 = .060
B) 0.40 + .15 = .55
C) 1 - (.40 + .15) = .45
D) 1 - (.6 * .85) = .490
Correct Answer:

Verified
Correct Answer:
Verified
Related Questions
Q32: The second set of branches of a
Q55: If the event of interest is A,the
Q76: In general,a posterior probability is calculated by
Q88: The _ of events A and B
Q124: Suppose the probability that a person owns
Q139: Messenger Service <br> Three messenger services deliver
Q145: An approach of assigning probabilities which assumes
Q148: The _ rule says that P(A<sup>c</sup>)= 1
Q152: Two events A and B are said
Q198: If A and B are any two